What is the density of a compound whose mass = 5.51001 g, and volume = 5.010 L? The formula for density is d = mass/volume



1
Expert's answer
2021-05-20T07:51:28-0400

Using the information given (the mass in grams and the volume in mililiters) we find the density of this substance as:


"d = \frac{m}{V} = \frac{5.51001\,g}{5.010\,L} = \frac{5.51001\,g}{5010\,mL}= 0.0010998\,g/mL"


In conclusion, the density of this substance is 1.0998 g/L

or 1.0998 X 10-3 g/mL